OpenStudy (kj4uts):

One advantage of polar equations is that you can use a polar function to describe a graph whose equation in rectangular coordinates would not be considered a function. True or False. I think the answer is true.

OpenStudy (kj4uts):

If the question said disadvantage I think that would be false. but since it is an advantage I think the answer is TRUE.

OpenStudy (perl):

hint, consider the polar equation r = 2. this is the graph of a circle with radius 2. but in rectangular equation it is x^2 + y^2 = 2^2 , which is more complicated

OpenStudy (perl):

it would be true

OpenStudy (perl):

my example illustrated that the polar equation is much easier to represent

OpenStudy (kj4uts):

oh ok

OpenStudy (perl):

r = 2 is simpler than x^2 + y^2 = 2^2 , right?

OpenStudy (kj4uts):

yes

OpenStudy (kj4uts):

@perl So because it is simpler than a rectangular function it makes it true?